how do i work out what X is if the Hypotenuse is 25, on side is x and the other is 2x????
Answer by mananth(16946)

one side is x
the other side = 2x
hypotenuse = 25
side1^2+side2^2 = hypotenuse^2
x^+(2x)^2=25^2
x^2+4x^2=625
5x^2=625
x^2=625/5
x^2=125
x=sqrt125
11.18 units